Direct Labor Variances

The following data relate to labor cost for production of 5,700 cellular telephones:

Actual: 3,880 hrs. at $13.1
Standard: 3,820 hrs. at $13.3

a. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Rate variance $   
Time variance $   
Total direct labor cost variance $   

b. The employees may have been less-experienced or poorly trained, thereby resulting in a   labor rate than planned. The lower level of experience or training may have resulted in   efficient performance. Thus, the actual time required was   than standard.
Answer #1

a. Rate variance = (standard rate - actual rate)*actual hours

= (13.3-13.1)*3880 = $776 favorable

Time variance = (3820-3880)*13.3 = 798 unfavorable

Total direct labor cost variance = 3820*13.3 - 3880*13.1 = $22 unfavorable

b. Actual time required was higher than standard.